OpenStudy (anonymous):

solve the system using the elimination method twice 7x - 4y = 10 6x + 3y = 8

OpenStudy (anonymous):

write the first equation in terms of x \[x= (4y +10)/7\] plug it into the x of the second equation \[6(4/7y + 10/7) + 3y = 8\] solve that equation for y and you will get \[y = -4/45\] plug that into the second equation and solve it for x \[6x + 3(-4/45) = 8\] you get \[x = 62/45 \] your answers are \[y = -4/45, x = 62/45\] if you plug these into both the equations they work